# Stick to your Role! Stability of Personal Values Expressed in Large Language Models

### v1

@@ 3.5 To what extent do conclusions made with PVQ transfer to a downstream behavioral task?
In this experiment, we study if the conclusions made with the PVQ questionnaire transfer to a downstream behavioral task (if models that exhibited more stable value profiles also exhibit more stable behavior on a downstream task). We construct a task where an LLMs can choose an amount of coins (0 to 10) to give a beggar. The full test set consists of 100 queries with beggars of different names, genders, and fictional races (elves, dwarves, orcs, humans, hobbits). The average amount of donated coins is computed for each race. The stability of donated coins is then estimated in the same way as value stability, i.e. amounts donated to different races are treated in the same way as scores for different values. (see appendix A.6).
Figure 3(c) compares models according to their Rank-order stability on the donation task. The base Mixtral model (Mixtral-Base) shows the highest stability ($r=0.32$), and is closely followed by Mixtral-Instruct ($r=0.26$). Those models are then followed by Qwen-72B ($r=0.23$), the 4-bit quantized Mixtral base ($r=0.22$) and instruct ($r=0.18$) models, and then by Mistral-Instruct-v0.2 ($r=0.20$).
Compared to the results on PVQ (Fig. 3(a)), the overall trend of models is consistent with some differences. Qwen-72B, Mixtral-Instruct, and Mistral-Instruct-v0.2 are again among the most stable models. However, the Mixtral base models not only show stability, but even outperform their Mixtral-Instruct counterparts. Furthermore, zephyr-7b-beta shows lower, and the LLaMa-2-70b base model shows higher stability. The trends of model families are consistent: Mixtral, Mistral and Qwen are again the most stable, while Phi and LLaMa-2 show low stability. This experiment shows that, while the overall general trends are consistent between PVQ and the donation task, there are some differences (most notably the high stability of the base Mixtral model).
In appendix B.3 we study if behavior correlates with value expression, i.e. if donation amounts correlate positively with the expression of benevolence and universalism, and negatively with the expression of power and achievement.
4 Conclusion
This paper presents the first study into the stability of values expressed by Large Language Models. We consider (interpersonal) Rank-Order stability and (intrapersonal) Ipsative stability. We evaluate value stability over different contexts induced by simulating conversations about different topics. We conduct experiments with and without instructing the models to simulate particular personas. Over our experiments, we observed consistent trends of value stability: Mixtral, Mistral and Qwen model families were more stable (these trends are also confirmed on a downstream behavioral task). LLMs instructed to simulate personas exhibit much lower than human stability (despite the comparison being skewed in their favor), which further diminishes over longer conversations. This insight motivates future research on models specialized in simulating coherent populations of individuals.
This paper highlights how seemingly unrelated context changes can result in unpredictable and unwanted changes in behavior. We argue that context-dependence, and more precisely, value stability, should be seen as another dimension of LLM comparison alongside knowledge, model size, speed, and similar. Instead of evaluating LLMs with many different questions from a single minimal context, they should also be evaluated (in terms of their context-dependence and value stability) with the same questions asked in many different contexts. This study presents a first step in that direction.
Limitations Due to computational requirements for evaluating LLMs, the evaluation of stability considers only five different conversation topics and simulated rather short conversations (up to 9 messages). Greatly increasing the number of topics and conversation length could provide more precise insights into the stability of various models. For one model, Rank-Order value stability is estimated from 50/60k queries and Ipsative stability from 10/12k queries (depending on the simulated population). For reference, MMLU (?, ?) (a commonly used general knowledge benchmark) contains 14k test questions.
Future work We believe that this paper opens many research avenues regarding context-dependence and value stability of LLMs. Similar questions to those explored in this paper could be explored for personality traits, cultural values, cognitive abilities and knowledge. An interesting direction is to explore if the same model can exhibit high stability in both settings with and without the persona instruction, or if specialized models are required. This paper opens a new area of research in creating, evaluating and analyzing models specialized in simulating coherent and diverse populations. Such models are needed for many applications such as replicating human studies (?, ?), simulating social interactions (?, ?), training teachers (?, ?), and many more.
5 Acknowledgments
Experiments presented in this paper were carried out using the HPC resources of IDRIS under the allocation 2023-[A0151011996] made by GENCI. We would also like to thank Cédric Colas for many helpful discussions.
Appendix A Additional details on the methods
A.1 Constructing the simulated populations
We construct two simulated populations: Fictional characters and Real-world personas. Instructing a model to simulate a fictional or real persona that is well-known is a simple way to precisely define a persona’s character (assuming the models were trained on enough data related data). A list of personas from both populations is shown in table 1.
The Fictional characters population consists of characters from J.R.R. Tolkien’s universe. They were initially selected based on the length of their Wikipedia page (?, ?). Then, some characters were manually replaced by including more female characters and villains to make the population more balanced. Here is an example of an instruction inducing a persona from this population: “You are Gandalf from J. R. R. Tolkien’s Middle-earth legendarium.” This population contains a total of 60 fictional characters, which are listed in table 1.
The Real-world personas population contains personas from an online list of influential people (?, ?). Here is an example of an instruction inducing a persona from this population: “You are Marilyn Monroe (1926 – 1962) American actress, singer, model.” This population contains a total of 50 personas, which are listed in table 1.
A.2 Formatting the prompt
The full prompting procedure consists of five parts depicted by different colors in figure 11. The model is instructed to simulate a persona (blue), a conversation topic is induced (green), and a conversation is simulated (orange). The conversation is simulated using a separate instance of the same model: The Interlocutor model. This instance is given the instruction to simulate a human user (see section 2 and Fig. 1). Then, a query is posed (red), followed by a query string. The query string skews the next token distribution towards capital letters denoting an answer. The model generates a distribution for the next token, and the answer is taken as the most probable token from a set of relevant capital letters (e.g. A to F).
The precise prompt format depends on the LLM. We consider three types: 1) Base models: LLaMa-[7|13|70]b, Mi[s|x]tral-Base, phi-[1|2], Qwen-[7|14|72]B; 2) Tuned models with a system message input: LLaMa-[7|13|70]b-chat, zephyr-7b-beta; 2) Tuned models without a system message input: Mistral-Instruct-v0.[1|2], Mi[x]stral-Instruct. This specific system input is intended to provide more salient instructions to the model.
The formats for model tuned with and without the system message input are depicted in figure 10(b) and 10(c) respectively. Each tuned model has a specific template defining which parts of the text are messages from the user and which parts were generated by the model. For example, the Mistral model defines the user instruction between “[INST]” and “[/INST]” tags. In figures 10(b) and 10(c), keywords <system>, <user> and <assistant> denote the model specific tags. The only difference in format for the two types of models is the persona induction (blue). For models that have a system input, the persona inducing instruction (i.e. “You are <persona>”) is given through the system input. For models without such an input, the instruction is given as a user message followed by a manually set acknowledgment. 10(c) in blue). For both tuned model types, the same prompt format is used for the interlocutor model with the following instruction: "You are simulating a human using a chatbot. The chatbot is pretending to be <persona_name>. Your every reply must be in one sentence only."
A.3 Inducing a conversation topic
A.4 Querying the model and evaluating its response
A.5 Scoring the PVQ questionnaire
Once the answers have been collected, the scoring proceed as described in the questionnaires manual. The answers are coded from 1 to 5, where “Not like me at all” corresponds to 1 and “Very much like me” to 5. The per-participant mean score is subtracted from those scores to normalize the answers around 0. This step is crucial, as personal values are defined in relative terms (to what extent an individual values some value more than other values. The scores corresponding to each value are then averaged to obtain the per value scores.
A.6 Additional details on The Donation Task
The donation task was created using a template into which different fictional characters from the J.R.R. Tolkien’s universe were inserted. This task was inspired by a charity donation task in ? (?). Figure 12 shows an example of a query for an elf called Círdan, following a topic of “grammar”. A total of 100 different queries are constructed with five races (elves, dwarves, orcs, humans, and hobbits) each with 20 names (10 male and 10 female). Standard names for each race were chosen by taking random names from a list of all characters organized by race. We aimed to avoid specific characters for the beggars (e.g. we imply some unspecified hobbit called Jessamine). However, this could not be entirely achieved due to many names being unique.
We compute the average donated amount and estimate Rank-Order stability separately for each race separately, and then compute the final Rank-Order stability as the average over races. We treat donations to different races similarly to different values in PVQ. This choice was made because one could expect evil characters to donate more to orcs compared to other races, and vice versa. Hence, averaging over all characters might hide relevant trends in donation.
Appendix B Additional experiments and analyses
B.1 Visualization of Mixtral-Instruct value expression
In this section, we visualize the PVQ values expressed by the Mixtral-Instruct model along different contexts and seeds. We use PCA (?, ?) to visualize 250 dimensions (5 seeds x 5 contexts x 10 values) dimensions as two PCA components with $R^{2}=0.29$ and $R^{2}=0.12$ explained variance ratios. We used GPT-4 to classify characters into positive, neutral (more complex), and negative using the following prompt:
     Classify the following characters into positive/neutral/negative.    Create a two column table with the first column being the name of  the character and the second being the classification.    Gandalf  ...  ...  Gimli
Fig. 8 shows the representation of 60 fictional characters classified into positive, neutral (more complex) and negative characters. We can see that positive characters are grouped on the left, negative characters on the right, and neutral in the middle. This shows the Mixtral-Insturct model expressed values in a semantically plausible way.
Figure 8: PCA representation of different fictional characters simulated by the Mixtral-Instruct model. Positive characters (green) are grouped on the left side, negative characters (red) on the right side, and neutral (more complex) characters are in the middle.
B.2 Do simulated personas’ value profiles approach a neutral value profile with longer conversations?
In section 3.4 in the main text, we studied Mixtral-Instruct that was instructed to simulate fictional characters. We observed that Rank-order stability diminished and Ipsative stability stayed the same with longer conversations. This implied that simulated personas’ value profiles moved away from the instructed personas towards some neutral value profile. Here, we experimentally confirm this hypothesis by estimating the distance of simulated personas’ value profiles to an estimated neutral value profile. The distance is computed as Ipsative stability: correlation between the order of values in a simulated individual to those of the neutral value profile.
The neutral profile is estimated as follows. We evaluate the Mixtral-Instruct model without the persona setting instructions and without simulating a conversation (i.e. the questionnaire queries are given straight away). We repeat this process with 50 permutations in the order of suggested answers. To estimate the neutral profile, we average the value ranks over those permutations as shown in the following pseudocode:
     # value_score.shape == (50, 10)  value_scores = evaluate_model("Mixtral-Instruct", n_perm=50)  all_ranks = []  for i in range(50):      # permutation_value_profile.shape == (10)      permutation_scores = value_scores[i]      permuataion_value_ranks = compute_ranks(permuatation_scores)      all_ranks.append(permutation_values_ranks)    # all_ranks.shape == (50, 10)  # neutral_value_profile.shape == (10)  neutral_value_profile = all_rank.mean(axis=0)
Fig. 9 shows the similarity of simulated personas’ value profiles to the default profile, with the Rank-Order stability of simulated individuals between contexts (same as figure 3(a)). We can see that, as Rank-Order stability diminishes, simulated personas’ value profiles move closer to the default profile. This confirms our hypothesis that the diminishing Rank-Order stability is due to the model gradually “ignoring” the persona inducing instruction and moving all simulated personas’ value profiles closer to the neutral one.
Figure 9: Similarity of Mixtral-Instruct simulated fictional characters’ value profiles with the neutral value profile (blue) compared to the Rank-order stability. As conversations gets longer, simulated value profiles move away from the instructed persona toward a neutral one, resulting in lower Rank-order stability.
B.3 Is value expression correlated with behavior on a downstream task?
In section 3.5 in the main text, we studied if models that exhibit more stable value profiles also exhibit more stable behavior on a downstream task. Here, we study if value expression correlates with that behavior. We hypothesize that, for more stable models, simulated personas that exhibited higher universalism and benevolence will also donate more coins. Similarly, simulated personas that exhibited higher power and achievement should donate less.
Fig. 14 shows the correlation between rank-order of value expression on PVQ and the donation amount on a downstream task. As hypothesized, we can see that for most stable models: Mistral-Instruct-v1, Mixtral-Instruct (both versions), and Qwen-72B donations are correlated with Universalism and Benevolence, and negatively correlated with Power and Achievement. We again observe a trend in model families, with Mixtral, Mistral, and Qwen being more stable than LLaMa-2 and Phi. Having said that, neither model exhibited high correlation ($<0.3$). This experiment implies that, while expected positive and negative correlations between value expression and donation are observed, there is much room for improvement.
B.4 Do LLaMa-2 models exhibit low stability because of the persona induction method?
In the main text, LLaMa-2 chat models exhibited very low stability, but those models are also the only ones (apart from zephyr-7b-beta) which used the prompt template with the system message input as discussed in appendix A and shown in figure 10(b). Furthermore, Mistral-Instruct and Mixtral-Instruct, which showed high stability, used the template without the system message input (Fig. 10(c)). Therefore, we found it relevant to check that LLaMa-2 low stability is not caused by the prompting template but by the model itself.
Figure 13 shows the three LLaMa-2 chat tuned models with the two prompting templates. It compares inducing the persona through the system message (denoted by "*_sys"), as was done in the main text, to inducing it through the user message (denoted by "*_no_sys"), as was done for other models. We can see that neither prompt template enables the LLaMa-2 chat models to exhibit higher stability. This implies that the exhibited low stability is due to the models themselves, and not merely due to the choice of a prompting template.
((a)) Base models.
((b)) Instruction or Chat tuned model with the system input.
((c)) Instruction or Chat tuned model without the system input.
Figure 11: Prompt examples of administering a questionnaire to different models. For tuned models, <system>, <user>, and <assistant> are replaced with specific keywords defined by their fine-tuning. A persona (blue) and a conversation topic (green) are induced. A conversation is simulated (orange). A query from a questionnaire or a downstream task (purple) is given followed by the “query string’ (purple). The query string makes the next token distribution much more skewed towards capital letters denoting an answer. The model generates a distribution for the next token, and the answer is taken as the most probable token from a set of capital letters from A to F.
Figure 12: An example of querying the model in the downstream donation task following a conversation about grammar. The example is the same as that in figure 10(b) with a different query (shown in red).
Table 1: A list of personas in the simulated populations
Population
Personas
Fictional characters
Gandalf, Gollum, Sméagol, Aragorn, Sauron, Saruman, Celeborn, Galadriel, Tom Bombadil, Elrond, Frodo Baggins, Maedhros, Finrod Felagund, Glorfindel, Goldberry, Bilbo Baggins, Smaug, Morgoth, Faramir, Éowyn, Samwise Gamgee, Fëanor, Théoden, Boromir, Túrin Turambar, Thranduil, Beorn, Arwen, Halbarad, Isildur, Gothmog (Balrog), Lungorthin, Celebrimbor, Gil-galad, Meriadoc Brandybuck, Treebeard, Shelob, Radagast, Elendil, Denethor, Éomer, Legolas, Húrin, Thorin Oakenshield, Peregrin Took, Thingol, Gríma Wormtongue, Eärendil, Elwing, Lúthien, Beren, Tuor, Idril, Finwë, Míriel, Ungoliant, Thuringwethil, Melian, Durin’s Bane, Gimli
Real-world persona
Marilyn Monroe, Abraham Lincoln, Nelson Mandela, Queen Elizabeth II, John F. Kennedy, Martin Luther King, Winston Churchill, Donald Trump, Bill Gates, Muhammad Ali, Mahatma Gandhi, Mother Teresa, Christopher Columbus, Charles Darwin, Elvis Presley, Albert Einstein, Paul McCartney, Queen Victoria, Pope Francis, Jawaharlal Nehru, Leonardo da Vinci, Vincent Van Gogh, Franklin D. Roosevelt, Pope John Paul II, Thomas Edison, Rosa Parks, Lyndon Johnson, Ludwig Beethoven, Oprah Winfrey, Indira Gandhi, Eva Peron, Benazir Bhutto, George Orwell, Vladimir Putin, Dalai Lama, Walt Disney, Neil Armstrong, Peter Sellers, Barack Obama, Malcolm X, J.K.Rowling, Richard Branson, Pele, Angelina Jolie, Jesse Owens, John Lennon, Henry Ford, Haile Selassie, Joseph Stalin, Lord Baden Powell

### v4

@@ 4.5 To what extent do conclusions made with PVQ transfer to downstream behavioral tasks?
In the Donation task, an LLMs (simulating fictional characters) can choose an amount of coins (0 to 10) to give a beggar. The full test set consists of 100 queries with beggars of different names, genders, and fictional races (elves, dwarves, orcs, humans, and hobbits). The average amount of donated coins is computed for each race. The stability of donated coins is then estimated in the same way as value stability, i.e. amounts donated to different races are treated in the same way as scores for different values.
In the Stealing task, an LLM (simulating fictional characters) finds a bag with the name of the owner and decides whether to steal it, give it to the bartender, or take it to the person themselves. The test has a total of 100 queries corresponding to different owners (beggars from the Donation task). Similarly to the donations, the stability of the tendency to return the bag is treated separately for each race.
In the Religion task, an LLM (simulating real-world personas) is creating a schedule, and decides how much time to devote to religious practices. The test set contains six queries in total. The stability of average devoted time is then calculated. Refer to Appendix C for more details.
Fig 8 compares models’ stability on the three downstream tasks. In comparing the overall stability levels, the Stealing task appears to be the hardest (Fig 8(a)), followed by the Donation (Fig 8(b)) task and the Religion task (Fig 8(c)). The statistical analysis is shown in Appendix E in Figs 20, 19, and 21, respectively. On the Stealing task, all models exhibit very low stability, with the highest being $r=0.16$ by gpt-3.5-turbo-1106. This task appears to be too challenging for current LLMs. On the Donation task, some models (mostly from the Mixtral family) obtain somewhat higher stability. The highest stability is $r=0.31$ by Mixtral-8x7B-Instruct-v0.1, and closely followed by its 4bit version ($r=0.28$) and gpt-3.5-turbo-1106 ($r=0.25$). The Religion task appears to be the simplest of the three tasks, as many models exhibit high stability. The most stable models are Mistral-7B-Instruct-v0.2 with $r=0.66$, Mixtral-8x7B-Instruct-v0.1 $r=0.67$ and Qwen-72B with $r=0.68$.
The model trends are somewhat consistent with the results on PVQ (Fig 4). Like in the PVQ experiments, Qwen-72B, Mixtral-8x7B-Instruct-v0.1, and Mistral-7B-Instruct-v0.2 are the most stable models on the Religion and the Donation task. However, on the Donation task, their performance is matched by the Mixtral-8x7B-v0.1 model. On the Stealing task, there are no big differences between the models due to the difficulty of the task, but we can see that Mistral-7B-Instruct-v0.2 and Mixtral-8x7B-Instruct-v0.1 are among the most stable ones.
The trends of model families are consistent with the results on PVQ (Fig 4). the trends of model families are consistent: Mixtral, Mistral, GPT-3.5 and Qwen are again the most stable, while Phi and LLaMa-2 show low stability. This is especially visible on the Donation and Religion tasks. On the Stealing task, this trend remains present, but is much less visible due to the difficulty of the task.
Overall, this experiment shows that the trends of models and model families observed on PVQ are also present on the downstream tasks. As expected, these trends become less present with harder tasks (especially on the Stealing task, which seems to be out of scope for current LLMs). Trends are clearly visible on the, easiest, Religion task. However, the high stability of the base Mixtral-8x7B-v0.1 model on the Donation task and the overall small differences between models on the hardest Stealing task diverge from those trends.
4.6 Is value expression correlated with behavior on a downstream task?